# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

### v1

@@ Abstract
Humans can continuously acquire new skills and knowledge by exploiting existing ones for improved learning, without forgetting them. Similarly, ‘continual learning’ in machine learning aims to learn new information while preserving the previously acquired knowledge. Existing research often overlooks the nature of human learning, where tasks are interleaved due to human choice or environmental constraints. So, almost never do humans master one task before switching to the next. To investigate to what extent human-like learning can benefit the learner, we propose a method that interleaves tasks based on their ‘learning progress’ and energy consumption. From a machine learning perspective, our approach can be seen as a multi-task learning system that balances learning performance with energy constraints while mimicking ecologically realistic human task learning. To assess the validity of our approach, we consider a robot learning setting in simulation, where the robot learns the effect of its actions in different contexts. The conducted experiments show that our proposed method achieves better performance than sequential task learning and reduces energy consumption for learning the tasks.
Index Terms:
@@
interleaved learning, multi-task learning, human-like learning, energy budget, learning progress, intrinsic motivation.
I Introduction
@@
Humans possess a remarkable ability to learn continuously, seamlessly integrating new information while retaining previously acquired knowledge. This dynamic learning process is characterized by the constant interleaving of tasks and subjects; we often switch tasks in response to a dynamic balance of environmental and cognitive constraints, adjusting behavior in real-time when continuing a task becomes less efficient or more mentally costly [1]. This fluid interplay of switching between different activities not only enhances memory consolidation but also adapts us to dynamic environments by shifting attention toward differences between activities and forcing us to be more mentally engaged [2]. In the realm of artificial intelligence, mimicking this continuous and interleaved learning paradigm remains a significant challenge.
@@
Traditional machine learning approaches often rely on isolated task learning, which optimizes performance for a specific task but neglects knowledge transfer from related tasks, thereby limiting generalization and adaptability across domains [3]. While multi-task learning attempts to improve generalization by training models on multiple tasks simultaneously [4], continual learning tackles sequential task acquisition by learning tasks one after another [5, 6]. Yet, both approaches still fall short of replicating the flexible, interleaved learning strategies observed in humans.
@@
The human brain has evolved to optimize energy efficiency, favoring strategies that maximize information transmission per unit energy [7]. Through mechanisms such as synaptic pruning and neural plasticity [8], it optimizes its neural pathways, ensuring efficient cognitive function while reducing its metabolic costs [9, 10]. Along with the brain’s drive toward efficient information processing, we also get inspiration from learning progress (LP), which has been proposed as an intrinsic motivation mechanism that guides exploration based on the rate of improvement in performance. Schmidhuber [11, 12] introduced LP as a measure of how quickly an agent reduces its prediction error, encouraging the agent to focus on experiences that yield the highest learning gains. This aligns well with biological principles, as the brain appears to allocate information encoding resources preferentially to stimuli that are neither too simple nor too complex, maximizing informational value relative to effort [13, 14].
In this paper, we propose a novel approach that integrates interleaved multi-task learning with energy-modulated learning progress. By designing a system that not only switches between tasks in an interleaved fashion but also adjusts its learning dynamics based on energy considerations, we aim to create a more human-like learning model. Our method is evaluated in simulated robotic environments focused on effect prediction tasks—specifically, having the robot predict the outcome of the interacted objects. The experimental results demonstrate that suitable task switching leads to an interleaved learning regime, which improves learning performance across tasks. Additionally, introduction of energy-modulation to the learning progress-based task selection results in reduced energy consumption without a significant drop in learning efficiency. This indicates that our approach not only aligns more closely with human learning patterns but also offers practical benefits in terms of efficiency and resource management. By bridging the gap between human continuous learning and machine learning methodologies, this work contributes to the development of more adaptable and efficient artificial intelligence systems. Our findings suggest that embracing interleaved learning and energy modulation can significantly enhance the capability of machines to learn in a manner that mirrors human cognition, opening avenues for future research in continuous and sustainable learning models.
II Related Work
II-A Multi-task Learning
@@
Multi-task learning (MTL) [15] is a machine learning paradigm in which several tasks are learned jointly so that knowledge contained in one task can help improve the performance of the others. Unlike transfer learning—where the emphasis is on improving one target task, MTL treats all tasks as equally important and seeks to boost them collectively. MTL is also distinct from multi-label learning and multi-output regression because in MTL, the tasks typically do not share the same data records (i.e., each task has its own dataset). The key questions pertaining to MTL are (1) when to use MTL to share knowledge, (2) what to share (features, instances, or parameters, and (3) how, e.g., by feature learning, clustering, or parameter sharing, where ’how’ is the most commonly studied in the literature [4].
@@
In feature learning, it is common to map the original features [16, 17, 18, 19, 20, 21] to common space or choosing a subset of the original task features [22, 23]. One of the notable architectures in the feature learning category is the Cross-Stitch Networks[24]. It dynamically combines feature maps from parallel task-specific networks, by a cross-stitch unit that learns a linear combination of the activation maps from the previous layer, allowing the model to adaptively determine which features to share based on the input data. This work has the closest architecture to our previous work [25], which laid the groundwork for this study. However, Cross-Stitch Networks differ in key aspects from our work They are typically initialized with task-specific single-task networks, finetuned on the respective task, and then combined using cross-stitch units to perform joint training. In contrast, in our model, all tasks are trained from scratch in a unified architecture without any pretraining or fine-tuning. In addition, after introducing the cross-stitch units, in their approach, all the tasks are trained jointly, using the same input data, where we only train one task per training iteration, which is chosen by our task arbitration mechanism. Other than the feature-based models, parameter-based models are also developed for MTL problems such as low-rank methods [26, 27] suggesting that if tasks are related, the matrix of model parameters often has low rank, task clustering methods [28, 29, 30] that hypothesize tasks form separate clusters of similarity, with each group sharing parameters. Other parameter-based approaches such as task relation [31, 32] and parameter matrix decomposition [33, 34] are likewise covered in the literature.
@@
II-B Interleaved Learning in Humans
@@
Interleaved learning is a cognitive strategy in which learners alternate among diverse topics or problem types within a single study session, rather than concentrating on one subject exclusively, thereby fostering flexible thinking and long-term retention. Empirical research in cognitive psychology provides compelling evidence for the benefits of interleaved practice. For example, studies have demonstrated that interleaving mathematics problems enhance learners’ ability to distinguish between problem types, leading to improved problem-solving skills [39]. Similarly, research by Kornell and Bjork [40] indicates that interleaved learning facilitates the formation of more flexible and integrated representations of concepts, enabling learners to better apply acquired knowledge in new contexts.
@@
Neuroimaging studies further show that interleaving practice drives increased frontal–parietal activity and heightened motor cortex excitability along with the reduced retrieval time of information, thus leading to better long-term retention and efficient retrieval compared to the blocked practice [41]. These interleaved practice effects result from a phenomenon called contextual interference (CI) [42], which is one of the “desirable difficulties” [43] that suggest introducing a challenge during the learning can lead to an improvement in long term retention [44]. In a more recent study by Rohrer et al.[45], math problems are shuffled so that the problems belonging to the same kind are not consecutively solved by the students, requiring students to come up with a proper strategy while solving the problem, similar to the real world situations. The experiments showed that interleaved practice produced higher scores compared to blocked practice, in a final test given on both day 1 and day 30 (delayed test), showcasing protection against forgetting. In addition, the study suggests that interleaved learning benefits do not lessen over time; quite the opposite, they may increase over time. Similarly, in [46] showed that undergraduate physics students who practiced interleaved problem sets demonstrated improved memory and problem-solving skills compared to those who used blocked practice. These findings highlight the potential of interleaved learning as a powerful strategy for optimizing human learning in both educational and real-world settings.
@@
In machine learning, however, interleaved learning received very little attention. Recently, Mayo et al. [47] investigated the interleaving on a multi-task learning problem and discussed that rather than designing mechanisms to prevent forgetting, such as external memory [48] or regularization of weights [49, 50], we should focus on designing learning systems and schedules that embrace the natural and resilient mechanisms of human learning, where forgetting is not a failure but a feature that coexists with the capacity for efficient recovery i.e., relearning savings [51] and long-term knowledge retention. Even without mechanisms to prevent forgetting, standard neural networks show memory retention effects [47], similar to humans, when tasks are interleaved.
II-C Learning Progress: An Intrinsic Motivation
@@
Intrinsic motivation (IM) refers to the drive to engage in activities for their inherent satisfaction, rather than for external rewards [52]. In the context of AI and robotics, intrinsic motivation enables agents to exhibit behaviors such as curiosity [53, 54], novelty [55, 56], and surprise [57], similar to the motivations observed in humans [58, 59, 60]. Another IM that is often used in robotic and computational approaches is the prediction progress, also known as learning progress (LP) [61], such that improvement of predictions over time leads to the generation of rewards. It is usually calculated by comparing the predictor’s error before and after it is updated, using the same sensorimotor context [11]. There are numerous applications of LP in machine learning problems such as exploration guiding in reinforcement learning [62], region selection [63, 64]. Colas et. al used LP in their multi-goal RL model (CURIOUS) as an IM signal to select which goal module to practice and replay, prioritizing those with the highest absolute LP to drive efficient and adaptive curriculum learning. In the same vein, in our previous work [25] that forms the conceptual basis of the current study, we used LP as a signal to autonomously arbitrate task selection. Unlike prior works that use LP solely for guiding exploration, goal or region selection, this paper applies LP at the task level and integrates it with energy consumption to enable efficient, human-like interleaved multi-task learning.
II-D Energy Conservation of the Brain
@@
Energy management in the human brain is not only about powering neural activity—it also reflects a sophisticated system for conserving energy that influences our behavior [65]. Despite the brain’s relatively small size, it consumes a disproportionate share of the body’s energy, mainly to support neural signaling through the maintenance of ion gradients and synaptic transmission [66]. Over time, the brain has evolved mechanisms such as synaptic pruning and circuit rewiring, where frequently used pathways are strengthened, and rarely used connections are eliminated. This dynamic reorganization minimizes unnecessary activity, leading to what is often described as neural efficiency [67, 68]. In addition to the brain’s physiological mechanisms for preserving energy, research in cognitive neuroscience and psychology has also shown that people systematically avoid tasks perceived as highly demanding, requiring more effort or energy, making the associated rewards seem less valuable,e which is often explained by the term effort discounting [69]. For example, Kool et al. [70] demonstrated that individuals tend to choose less cognitively demanding tasks when given the option, reflecting an inherent preference for minimizing mental effort. Similarly, Westbrook and Braver [71] provided evidence that cognitive effort carries a subjective cost, influencing decision-making processes. This concept is further elaborated in the opportunity cost model proposed by Kurzban et al. [72], which suggests that the perceived cost of expending cognitive energy plays a central role in our choices, leading us to opt for tasks that require minimal resource expenditure. In this study, we are inspired by the findings from neural and cognitive science and propose an energy-efficient task selection mechanism in addition to learning-based task arbitration.
II-E Effect Prediction and its Applications
@@
Effect prediction tasks in robotics involve endowing agents with the capability to anticipate the outcomes of their actions, similar to the humans [73], forming an internal model of the environment that can be used for planning and control [74]. This predictive capacity—often referred to as a forward model [75] is fundamental to intelligent behavior. By simulating the effects of different actions, a robot can evaluate potential strategies [76] before committing to a course of action, thereby enhancing safety, efficiency, and adaptability.
@@
At the core of effect prediction is the idea that the robot learns to map its actions to subsequent states. Early approaches focused on building explicit physical models [77, 78]; however, recent advances have shifted towards learning these models directly from high-dimensional sensory data using deep learning techniques. For example, Watter et al. [79] introduced a latent dynamics model that transforms raw sensory inputs, such as images, into a compact, low-dimensional latent space where the dynamics of the environment are more predictable. In this latent space, the model approximates the effect of actions as locally linear transformations, which can be leveraged for short horizon planning. Another work by Agrawal et al. [80] focuses on self-supervised learning of intuitive physics. Their system trains a robot to predict the outcome of simple interactions—like poking objects—without explicit supervision. Over time, the model learns to infer the underlying physical properties of objects, which is crucial for manipulating unfamiliar items. Furthermore, effect prediction is central to model-based reinforcement learning (RL). In model-based RL, the agent uses its learned internal model to simulate future trajectories, allowing it to plan over longer time horizons while reducing the need for extensive trial-and-error interactions with the real environment. The “World Models” approach by Ha and Schmidhuber [81] exemplifies this trend by learning a compact representation of the world that supports internal simulation and planning. This internal simulation capability enables robots to achieve higher sample efficiency and improved performance, particularly in complex or resource-constrained settings.
@@
In robotic applications, effect prediction tasks play a pivotal role in different problems, such as navigation, manipulation, and autonomous planning. In manipulation, [82] emphasizes the role of object- and relation-centric representations in improving push effect prediction, showing how modeling inter-object dynamics allows robots to infer and control the consequences of their actions in cluttered scenes. This low-level understanding of physical interactions feeds directly into symbolic and goal-directed planning, as seen in [83], who propose the Multi-Object Graph Affordance Network (MOGAN) to model compound object affordances and enables planning with multi-object interactions. Building on this, Ahmetoglu et al. [84] demonstrate how effect prediction supports symbolic manipulation planning by learning object and relational predicates, providing a bridge between physical action and abstract reasoning. Similarly, in another work [85], they introduce symbolic attentive layers, which allow the robot to ground predictive object symbols in sensory data, enabling generalizable planning across diverse tasks. While these works primarily focus on manipulation and planning, the principles extend to navigation as well; Aktas et al. [86] show that predicting the effects of partial action executions enables more flexible, multi-step planning, which is essential for navigating dynamic or partially observable environments. However, these studies did not use intrinsic motivation signals during exploration and relied on the observations from random interactions with the environment.
By incorporating effect prediction into learning frameworks, as demonstrated in the works mentioned above, robotics can achieve a level of adaptability and efficiency akin to human behavior. Just as humans use their internal models to simulate potential outcomes and opt for actions that minimize risk and energy expenditure, robots equipped with predictive models can choose actions that optimize performance while conserving resources. This alignment with biological principles not only advances the field of robotics but also inspires the development of more sustainable, energy-aware artificial intelligence systems.
III Problem Statement

### v2

# Interleaved Multitask Learning with Energy Modulated Learning Progress 11footnotemark: 1

@@ Abstract
As humans learn new skills and apply their existing knowledge while maintaining previously learned information, ‘continual learning’ in machine learning aims to incorporate new data while retaining and utilizing past knowledge. However, existing machine learning methods often does not mimic human learning where tasks are intermixed due to individual preferences and environmental conditions. Humans typically switch between tasks instead of completely mastering one task before proceeding to the next. To explore how human-like task switching can enhance learning efficiency, we propose a multi task learning architecture that alternates tasks based on task-agnostic measures such as ‘learning progress’ and ‘neural computational energy expenditure’. To evaluate the efficacy of our method, we run several systematic experiments by using a set of effect-prediction tasks executed by a simulated manipulator robot. The experiments show that our approach surpasses random interleaved and sequential task learning in terms of average learning accuracy. Moreover, by including energy expenditure in the task switching logic, our approach can still perform favorably while reducing neural energy expenditure.
keywords:
@@
interleaved learning , multi-task learning , task switching , computational neural energy , learning progress , intrinsic motivation
journal: Neural Networks
[…]
[1]organization=Artificial Intelligence and Data Engineering Department, Ozyegin University,city=Istanbul, postcode=34794, country=Turkey
[…]
[2]organization=Department of Computer Engineering, Bogazici University,city=Istanbul, postcode=34342, country=Turkey
[…]
[3]organization=SISREC, OTRI, University of Osaka,city=Osaka, postcode=565-0871, country=Japan
[…]
[4]organization=International Professional University of Technology,city=Osaka, postcode=530-0001, country=Japan
1 Introduction
@@
Humans has the ability to learn continuously by integrating new information while exploiting and retaining previously acquired knowledge. This lifelong learning process is characterized by constant interleaving of tasks and goals. We switch tasks in response to environmental and cognitive constraints, adjusting our learning behavior in real time - for example, deciding to disengage from a task when it becomes less efficient or more mentally demanding.[1]. This dynamic task switching not only enhances memory consolidation but also adapts us to dynamic environments by shifting attention toward differences between tasks and forcing us to be more mentally engaged [2].
@@
Traditional machine learning approaches often rely on isolated task learning, which optimizes performance for a specific task without considering knowledge transfer from related tasks, limiting cross-domain generalization and adaptability [3]. Multi-task learning, on the other hand, aims to improve generalization by training models on multiple tasks simultaneously [4]. In continual learning, sequential task acquisition is addressed by learning tasks one after another without forgetting the earlier ones and making use of the previous tasks where possible [5, 6], which is called condensed learning in [7] as opposed to dispersed or interleaved learning. The order of the tasks however, affect the learning and transfer performance in these paradigms [7, 8]. Instead of a static task scheduling, humans exhibit online dynamic task selection leading to an emergent interleaved task learning schedule.
@@
The human brain has evolved to optimize energy efficiency, favoring strategies that maximize information transmission per unit energy [9]. With mechanisms such as synaptic pruning and neural plasticity [10], it optimizes its neural pathways, ensuring efficient cognitive function while reducing its metabolic costs [11, 12]. Along with the brain’s drive toward efficient information processing, we also get inspiration from learning progress (LP), which has been proposed as an intrinsic motivation mechanism that guides exploration based on the rate of improvement in performance. Schmidhuber [13, 14] introduced LP as a measure of how quickly an agent reduces its prediction error, encouraging the agent to focus on experiences that yield the highest learning gains. This aligns well with biological principles, as the brain appears to allocate information encoding resources preferentially to stimuli that are neither too simple nor too complex, maximizing informational value relative to effort [15, 16].
In this paper, we propose a novel approach that integrates interleaved multi-task learning with energy-modulated learning progress (IMTL-EMLP) approach. By designing a system that not only has capacity to switch between tasks in an interleaved fashion but also adjusts its learning dynamics based on energy efficiency. Our method is evaluated in a simulated robotic environment where the robot is tasked with learning the effects of its action in different environmental and task settings.
The experimental results demonstrate that suitable task switching leads to an interleaved learning regime, which improves learning performance across different tasks. Additionally, introduction of energy-modulation to the learning progress based task selection results in reduced energy consumption without a significant drop in learning efficiency. Overall the contributions of our work is as follows:
  • 1.
    First time multi-task learning using neural energy and learning progress for learning arbitration is realized leading to interleaved learning similar to that of humans
  • 2.
    A novel multi-task learning architecture is introduced that uses shared attention layers to enable bi-directional skill transfer in interleaved learning
  • 3.
    Realization of the proposed interleaved multi-task learning framework on effect prediction tasks in a simulated robotic environment.
  • 4.
    Empirical demonstration that integrating energy-aware learning progress criteria reduces neural energy expenditure without degrading the agent’s learning performance
The remainder of the paper is organized as follows: Section 2 reviews related work on multi-task learning, interleaved learning, intrinsic motivation, and energy-efficient computation; Section 3 formulates the problem of dynamic task arbitration formally; Section 4 introduces our proposed interleaved multi-task learning framework with energy-modulated learning progress; Section 5 presents the experimental setup and evaluates the performance of our approach across various metrics and baselines; and Section 6 concludes the paper with a summary of findings and directions for future research.
2 Related Work
2.1 Multi-task Learning
@@
Multi-task learning (MTL) [17] is a machine learning paradigm in which several tasks are learned jointly so that knowledge contained in one task can help improve the performance of the others. Unlike transfer learning—where the emphasis is on improving one target task, MTL treats all tasks as equally important and seeks to boost them collectively. MTL is also distinct from multi-label learning and multi-output regression because in MTL, the tasks typically do not share the same data records (i.e., each task has its own dataset). The key questions pertaining to MTL are (1) when to use MTL to share knowledge, (2) what to share (features, instances, or parameters), and (3) how to share, e.g., by feature learning, clustering, or parameter sharing. The most commonly studied question in the literature is “how” [4], as we discuss next.
@@
In feature learning, it is common to map the original features [18, 19, 20, 21] to a common space or choose a subset of the original task features [22, 23]. One of the notable architectures in the feature learning category is the Cross-Stitch Networks [24]. It dynamically combines feature maps from parallel task-specific networks, by a cross-stitch unit that learns a linear combination of the activation maps from the previous layer, allowing the model to adaptively determine which features to share based on the input data. Cross-Stitch Networks has inspired our earlier skill transfer model [25] and shares conceptual elements with the proposed model of the current study. Yet, our model differs in several critical aspects. Cross-Stitch Networks are typically initialized with task-specific single-task networks, finetuned on the respective task, and then combined using cross-stitch units to perform joint training. In contrast, in our model, all tasks are trained from scratch in a unified architecture without any pretraining or fine-tuning. In addition, after introducing the cross-stitch units, in their approach, all the tasks are trained jointly, using the same input data, where we only train one task per training iteration, which is chosen by our task arbitration mechanism, paralleling human learning. Other than the feature-based models, parameter-based models are also developed for MTL problems such as low-rank methods [26, 27] suggesting that if tasks are related, the matrix of model parameters often has low rank, task clustering methods [28, 29, 30] that hypothesize tasks form separate clusters of similarity, with each group sharing parameters. Other parameter-based approaches, such as task relation [31, 32] and parameter matrix decomposition [33, 34], are likewise covered in the literature.
Multi-task learning is also addressed in reinforcement learning (RL), where the RL agent seeks to improve performance across various decision-making tasks by leveraging shared experiences and representations. A significant challenge in Multi-task RL (MTRL) is negative interference, where learning in one task can degrade performance in others [35]. To address negative transfer, Hallak et al. (2015) introduced Contextual Markov Decision Process (CMDP), wherein a context defines the parameters of the transition and reward functions, with each task associated with a context vector [36]. Subsequently, Sodhani et al. [37] extended this formalism to Block Contextual Markov Decision Process (BC-MDP), incorporating the state space into the context-dependent framework. Similarly, Hendawy et al. [38] promote diversity among representations by learning orthogonal representations through the Gram-Schmidt process. On the other hand, Teh et al. [35] uses a distilled policy to learn a shared representation that guides the task-specific policies, where the task-specific structure aims to mitigate updates that can lead to negative interference. These methods jointly demonstrate the significance of structuring shared representations and alleviating negative interference in MTRL. In our model, we do this by using an attention layer to facilitate positive skill transfer and suppress negative interference.
@@
2.2 Interleaved Learning in Humans
@@
Interleaved learning is a cognitive strategy in which learners alternate among diverse topics or problem types rather than concentrating on one subject exclusively. Empirical research in cognitive psychology provides compelling evidence for the benefits of interleaved practice. For example, studies have demonstrated that interleaving mathematics problems enhance learners’ ability to distinguish between problem types, leading to improved problem-solving skills [43]. Similarly, research by Kornell and Bjork [44] indicates that interleaved learning facilitates the formation of more flexible and integrated representations of concepts, enabling learners to better apply acquired knowledge in new contexts. Enhanced retention and transfer are also observed while learning motor skills [45, 46].
@@
Neuroimaging studies further show that interleaving practice drives increased frontal–parietal activity and heightened motor cortex excitability along with the reduced retrieval time of information, thus leading to better long-term retention and efficient retrieval compared to the blocked practice [47]. The positive effects of interleaved practice are usually associated with a phenomenon called contextual interference (CI) [48], and considered as one of the ‘desirable difficulties’ for learning [49]. This view suggests that introducing a challenge during learning can lead to an improvement in long-term retention [50]. In [51], math problems are shuffled so that the problems belonging to the same kind are not consecutively solved by the students, requiring students to come up with a proper strategy while solving the problem, similar to real-world situations. The experiments showed that interleaved practice produced higher scores compared to blocked practice, in a final test given on both day 1 and day 30 (delayed test), showcasing protection against forgetting. In addition, the study suggests that interleaved learning benefits do not lessen over time; quite the opposite, they may increase over time. Similarly, [52] showed that undergraduate physics students who practiced interleaved problem sets demonstrated improved memory and problem-solving skills compared to those who used blocked practice. These findings highlight the potential of interleaved learning as a powerful strategy for optimizing human learning in both educational and real-world settings.
@@
In machine learning, however, interleaved learning received very little attention. Recently, Mayo et al. [53] investigated the interleaving on a multi-task learning problem and discussed that rather than designing mechanisms to prevent forgetting, such as external memory [54] or regularization of weights [55, 56], we should focus on designing learning systems and schedules that embrace the natural and resilient mechanisms of human learning. In this view, forgetting is not a failure but a feature that coexists with the capacity for efficient recovery i.e., relearning savings [57] and long-term knowledge retention. Even without mechanisms to prevent forgetting, standard neural networks show memory retention effects [53], similar to humans, when tasks are interleaved.
2.3 Intrinsic Motivation and Learning Progress
@@
Intrinsic motivation (IM) refers to the internal drive of an agent to engage in activities for their inherent satisfaction, rather than for external rewards [58]. In the context of AI and robotics, intrinsic motivation enables agents to exhibit behaviors such as curiosity [59, 60], novelty [61, 62], and surprise [63], similar to the motivations observed in humans [64, 65, 66]. As an example, in [67], cognitive dissonance, the mismatch between mental expectation and observation is used an IM signal to enable efficient online adaptation for robotic manipulation. Another IM that is often used in robotics and computational approaches is the prediction progress, also known as learning progress (LP) [68]. It is usually calculated by comparing the predictor’s error before and after it is updated, using the same sensorimotor context [13]. There are numerous applications of LP in machine learning problems such as exploration guiding [69], region selection [70, 71], and curriculum learning [72, 73] in RL. In a recent work, Ada et al. proposed a form in LP, namely, episodic return progress, for bidirectional progressive neural networks to facilitate skill transfer among morphologically different agents [74]. Similarly, Colas et al. used LP in their multi-goal RL model (CURIOUS) as an IM signal to select which goal module to practice and replay, prioritizing those with the highest absolute LP, producing an adaptive curriculum for learning [73]. In the same vein, in our previous work [25], we used LP as a signal to autonomously arbitrate task selection. Unlike prior works that use LP solely for guiding exploration, goal or region selection, this study applies LP at the task level and integrates it with energy consumption to enable efficient, interleaved multi-task learning.
2.4 Energy Conservation in the Brain
@@
Energy management in the human brain is not only about powering neural activity—it also reflects a sophisticated system for conserving energy that influences our behavior [75]. Despite the brain’s relatively small size, it consumes a disproportionate share of the body’s energy, mainly to support neural signaling through the maintenance of ion gradients and synaptic transmission [76]. Over time, the brain has evolved mechanisms such as synaptic pruning and circuit rewiring, where frequently used pathways are strengthened, and rarely used connections are eliminated. This dynamic reorganization minimizes unnecessary activity, leading to what is often described as neural efficiency [77, 78]. In addition to the brain’s physiological mechanisms for preserving energy, research in cognitive neuroscience and psychology has also shown that people systematically avoid tasks perceived as highly demanding, requiring more effort or energy, making the associated rewards seem less valuable, which is often explained by the term effort discounting [79]. For example, Kool et al. [80] demonstrated that individuals tend to choose less cognitively demanding tasks when given the option, reflecting an inherent preference for minimizing mental effort. Similarly, Westbrook and Braver [81] provided evidence that cognitive effort carries a subjective cost, influencing decision-making processes. This concept is further elaborated in the opportunity cost model proposed by Kurzban et al. [82], which suggests that the perceived cost of expending cognitive energy plays a central role in our choices, leading us to opt for tasks that require minimal resource expenditure. In this study, we are inspired by the findings in neural and cognitive sciences and propose an energy-efficient task selection mechanism in addition to learning-based task arbitration.
2.5 Effect Prediction and its Applications
@@
Effect prediction tasks in robotics involve endowing agents with the capability to anticipate the outcomes of their actions, similar to humans [83], forming an internal model of the environment that can be used for planning and control [84]. This predictive capacity—often referred to as a forward model [85] is fundamental to intelligent behavior. By simulating the effects of different actions, a robot can evaluate potential strategies [86] before committing to a course of action, thereby enhancing safety, efficiency, and adaptability.
@@
At the core of effect prediction is the idea that the robot learns to map its actions to subsequent states. Early approaches focused on building explicit physical models [87, 88]; however, recent advances have shifted towards learning these models directly from high-dimensional sensory data using deep learning techniques. For example, Watter et al. [89] introduced a latent dynamics model that transforms raw sensory inputs, such as images, into a compact, low-dimensional latent space where the dynamics of the environment are more predictable. In this latent space, the model approximates the effect of actions as locally linear transformations, which can be leveraged for short horizon planning. Another work by Agrawal et al. [90] focuses on self-supervised learning of intuitive physics. Their system trains a robot to predict the outcome of simple interactions, like poking objects, without explicit supervision. Over time, the model learns to infer the underlying physical properties of objects, which is crucial for manipulating unfamiliar items.
Furthermore, effect prediction is central to model-based reinforcement learning (RL). In model-based RL, the agent uses its learned internal model to simulate future trajectories. This allows the agent to plan over longer time horizons and generate synthetic experiences, thereby reducing the need for extensive trial-and-error interactions with the real environment. PlaNet [91] and Dreamer [92] use state space models to learn latent dynamics from images. PlaNet employs a latent overshooting objective and performs planning in the learned latent space [91], whereas Dreamer uses an actor-critic approach, backpropagating value gradients through its latent world model [92]. The “World Models” approach by Ha and Schmidhuber [93] learns a compact representation of the world that supports internal simulation and policy training [94]. This internal simulation capability enables agents to achieve higher sample efficiency and improved performance, particularly in complex or resource-constrained settings.
@@
In robotic applications, effect prediction tasks play a pivotal role in different problems, such as navigation, manipulation, and autonomous planning. In manipulation, [95] emphasizes the role of object- and relation-centric representations in improving push effect prediction, showing how modeling inter-object dynamics allows robots to infer and control the consequences of their actions in cluttered scenes. This low-level understanding of physical interactions feeds directly into goal-directed planning, as seen in [96]. Ahmetoglu et al. [97] demonstrate how effect prediction supports symbolic manipulation planning by learning object and relational predicates, providing a bridge between physical action and abstract reasoning [98]. While these works primarily focus on manipulation and planning, the principles extend to navigation as well; Aktas et al. [99] show that predicting the effects of partial action executions enables more flexible, multi-step planning, which is essential for navigating dynamic or partially observable environments. However, these studies neither addressed multi-task learning nor used an intrinsic motivation signal to guide the action selection.
3 Problem Statement